OpenStudy (anonymous):

Renting a pavilion for a birthday party costs $100 for the first hour and $75 for each additional hour. Show your work.

OpenStudy (anonymous):

What given values? What have you done so far?

OpenStudy (campbell_st):

if I write it this way can you identify the common factor \[5x \times 1 \] \[10x^2 = 5x \times 2x\] \[20x^3 = 5x \times 4x^2\] hope that helps